OpenStudy (anonymous):

What is the Laplace transform of 2t^5 + 4t^2 + 3t + 2???

OpenStudy (lgbasallote):

you can break it down right? \[\mathcal L \{2t^5\} + \mathcal L \{4t^2\} + \mathcal L \{3t\} + \mathcal L\{2\}\]

OpenStudy (lgbasallote):

then you can just take out the constants \[2\mathcal L \{t^5\} + 4\mathcal L \{t^2\} + 3\mathcal L \{t\} + 2 \mathcal L\{1\}\]

OpenStudy (anonymous):

That's what i had, just wanted to check i had done it right! Haven't got any answers for this paper :) Thank You

OpenStudy (lgbasallote):

welcome ^_^ so i assume you know how to proceed from there right?

OpenStudy (anonymous):

I've got \[240/s ^{6} + 24/s ^{3} + 6/s ^{2} + 2/s\] Does that look right? So bad at maths!

OpenStudy (lgbasallote):

the second term...why 24?

OpenStudy (lgbasallote):

that's \[\huge 4 \mathcal L \{t^2\}\] right? or did i write the exponent wrong?

OpenStudy (lgbasallote):

the third term too..

OpenStudy (lgbasallote):

it should be \[\Large` \mathcal L \{2t^5 + 4t^2+3t+2\} = \frac{240}{s^6} + \frac{8}{s^3} + \frac{3}{s^2} + \frac 2s\]

OpenStudy (anonymous):

Thank You!!!! Laplace is doing my head in - you've been a huge help!

OpenStudy (lgbasallote):

welcoem ^_^